OpenStudy (amtran_bus):

Please help me easy algebra f(x)= 4x^2-x+4 Find 2f(a) and [f(a)]^2

OpenStudy (amtran_bus):

@jim_thompson5910

zepdrix (zepdrix):

You're a bus +_+

OpenStudy (amtran_bus):

I like AmTran buses, I am one!

zepdrix (zepdrix):

\[\Large\rm f(\color{royalblue}{x})=4\color{royalblue}{x}^2-\color{royalblue}{x}+4\]If we multiply by 2,\[\Large\rm 2f(\color{royalblue}{x})=2(4\color{royalblue}{x}^2-\color{royalblue}{x}+4)\]And then plug a in for our x,\[\Large\rm 2f(\color{royalblue}{a})=2(4\color{royalblue}{a}^2-\color{royalblue}{a}+4)\]

OpenStudy (amtran_bus):

Thank you! Makes sense!

zepdrix (zepdrix):

So then uhhhh, how bout the next one? Shouldn't be too bad, right? Plug in a, then square each side.

OpenStudy (amtran_bus):

The [()] messes me up.

OpenStudy (amtran_bus):

@zepdrix

zepdrix (zepdrix):

They just added another set of brackets because \(\Large\rm f(a)^2\) looks really sloppy.

zepdrix (zepdrix):

There is no difference between square and round brackets. We just use different brackets to make it easier to match them up sometimes.

OpenStudy (amtran_bus):

Thank you.

zepdrix (zepdrix):

\[\Large\rm f(\color{royalblue}{x})=4\color{royalblue}{x}^2-\color{royalblue}{x}+4\]So plugging in a gives you,\[\Large\rm f(\color{royalblue}{a})=4\color{royalblue}{a}^2-\color{royalblue}{a}+4\]Squaring each side gives you,\[\Large\rm \left(f(\color{royalblue}{a})\right)^2=\left(4\color{royalblue}{a}^2-\color{royalblue}{a}+4\right)^2\]See how the double round brackets are a little harder to read?

zepdrix (zepdrix):

From there, your teacher might want you to expand out the brackets.. I'm not sure :o

OpenStudy (amtran_bus):

Can you help me find this domain? Find domain of y=(1/2x)-(11/2)

OpenStudy (amtran_bus):

Find domain of y=(1/2x)-(11/2)

OpenStudy (amtran_bus):

Can you help me with that? Thanks for working that other out.